Jramirez:

Henry and Wayne operate a charter flying service out of Breckenridge, which sits at 9,600 feet above sea level. Henry has two flights scheduled for the day. He is taking a couple from Breckenridge to Hamilton with a drop of 35% in elevation. At Hamilton, he is picking up 3 business men and taking them to Denver with a drop in elevation of 15%, which puts them at an elevation of feet above sea level. Wayne also has two flights scheduled for the day. He is taking a family from Breckenridge to Colorado Springs with a drop in elevation. At Colorado Springs, he is picking up some fishermen and taking them to Boulder with a drop in elevation, which puts them at an elevation of feet above sea level. After Henry and Wayne have both finished their two flights of the day, the final difference in their elevations is feet.

Gdeinward:

Heres the full directions: Directions: Type the correct answer in each box. Use numerals instead of words. Henry and Wayne operate a charter flying service out of Breckenridge, which sits at 9,600 feet above sea level. Henry has two flights scheduled for the day. He is taking a couple from Breckenridge to Hamilton with drop of 35% in elevation. At Hamilton, he is picking up 3 business men and taking them to Denver with a drop in elevation of 15%, which puts them at an elevation of_____feet above sea level. Wayne also has two flights scheduled for the day. He is taking a family from Breckenridge to Colorado Springs with a 3/8 drop in elevation. At Colorado Springs, he is picking up some fishermen and taking them to Boulder with a 1/10 drop in elevation, which puts them at an elevation of feet above sea level. After Henry and Wayne have both finished their two flights of the day, the final difference in their elevations is______feet. H=9,600-35% - 15% H=9,600- 50% H= 9600 * 0.50 Henry's elevation is 9600 * 0.50 W = 9,600 - 3/8 - 1/10 W = (9,600 * 0.375) * 0.1 Waynes elevation is (9,600 * 0.375) * 0.1 Henries elevation - Waynes elevation = answer 9600 * 0.50 - (9,600 * 0.375) * 0.1 = answer Do you understand the answer? The only reason I didnt finish it is because you must do some of the work. We cannot give direct answers.